A 5.0 kg box is being dragged with a rope. Find the magnitude of the tension necessary to keep the box moving at a constant velocity.


1
Expert's answer
2022-03-10T11:15:32-0500

The Newton's second law says

"ma=F_{net}=T-mg"

When the box moves with constant velocity the acceleration is zero. So

"T=mg\\
=5.0*9.8=49\:\rm N"